# Quantifying artificial intelligence through algebraic generalization

### v1

@@ Conclusion
Quantifying the algebraic ability of AI systems is difficult due to the lack of a theoretical framework from which to establish meaningful benchmarks. While there has been an increasing number of studies that have employed algebraic and compositional tasks to reliably elicit failure modes of transformers and LLMs, no mathematical framework exists to interpret these findings. Here, we provide a parsimonious framework – algebraic circuit complexity – to evaluate the extent of a model’s algebraic generalization ability in terms of their circuit divergences. In contrast to other formulations of complexity, such as Kolmogorov Complexity, encoding algebraic problems as circuits provides an explicitly computable formulation. The rich expressivity of algebraic problems, the data-rich nature of producing algebraic datasets, and the close links with algebraic circuits to other mathematical fields, makes algebraic circuit complexity a fruitful approach to quantify symbolic generalization in AI systems. We hope this lays the theoretical groundwork for future studies to quantify algebraic generalization in modern AI systems.
Acknowledgements
We thank Marco Carmosino for helpful discussions on circuit complexity and symbolic generalization. We acknowledge funding support from the Exploratory Science Councils at IBM Research.

### v2

@@ Conclusion
Quantifying the algorithmic ability of AI systems is difficult due to the lack of a theoretical framework from which to establish meaningful benchmarks. While there has been an increasing number of studies that have employed algebraic and compositional tasks to reliably elicit failure modes of transformers and LLMs, no theoretical framework exists to interpret these findings. In this Perspective, we provide a parsimonious framework – algebraic circuit complexity – to evaluate the extent of a model’s algorithmic generalization ability in terms of their circuit divergences. In contrast to other formulations of complexity, such as Kolmogorov Complexity, encoding algorithmic problems as circuits provides an explicitly computable formulation. The rich expressivity of algebraic problems, the data-rich nature of producing algebraic datasets, and the close links with algebraic circuits to other mathematical fields, makes algebraic circuit complexity a fruitful approach to quantify algorithmic generalization in AI systems. More generally, the circuit complexity framework introduced here (with algebraic functions) can be naturally extended to other computational problems (e.g., formal languages, Boolean circuits). We hope this Perspective provides the theoretical groundwork for future studies to quantify algorithmic generalization in modern AI systems with circuits.
Acknowledgments
We thank Marco Carmosino and Karan Srivastava for helpful discussions on earlier versions of the manuscript. We acknowledge funding support from the Exploratory Science Councils at IBM Research.
Competing interests statement
The authors declare no competing interests.